Taking with example write the mechanism of Mannich reaction?
1
Expert's answer
2020-04-06T12:00:01-0400

CONVERSION OF "HCHO"

"HCHO+NHR_2\to O^{-}CH_2-N^{+}HR_2\to OH-CH_2-NR_2"

"OH-CH_2-NR_2+H^{+}\to CH_2=N^{+}R_2+H_2O"

"CH_2=N^{+}R_2\to C^{+}H_2-NR_2"

CONVERSION OF "R-CH_2-CRO"

"R-CH_2-CRO\to R-CH=CRO^{-}+H^{+}"

FINAL STEP

"R-CH=CRO^{-}+CH_2-N^{+}R_2\to NR_2-CH_2-CHR-CRO"